The beams and pans of a balance have negligible mass. An object weighs W1 when placed in one pan and W2 when placed in the other . The weight W of the object is

Grade:

1 Answers

sahil agrawal
4 years ago
Let the weight of mass used for measuring be W. so by torque equilibrium W×x=W1×y. and second time W×y=W2×x.. on multiplying W1×W2=W^2. so W=√W1W2